OpenStudy (anonymous):

Show all work. Find the quotient: y^4 + 2y-3 /y^2+8

OpenStudy (anonymous):

Are you missing parentheses? Is that (y^4+2y-3)/(y^2+8)?

OpenStudy (anonymous):

Sure, it is written out like a fraction, so yes it probably needs parentheses.

OpenStudy (anonymous):

Its a different One ..but maybe i can get it... :)

OpenStudy (anonymous):

I would help work it out but I am awful with division

OpenStudy (anonymous):

y^2 -8 ------------------------------------------- y^2 +8 | y ^4 +2y -3 | | - (y^4 + 8y^2) ------------------- -8y^2 +2y -3 -( -8y^2 -64) -------------- 2y + 61 So maybe y^2 -8 is the quotient.........And 2y +61 remainder

OpenStudy (anonymous):

I don't Know whether its correct or not......just did it.. :) If its wrong then pls do tell me..

OpenStudy (anonymous):

Nickie_S Pls tell me am i correct?

OpenStudy (anonymous):

I wouldn't know, I am awful at this stuff, its good enough for me. lol:)

OpenStudy (anonymous):

U must have the ans..

OpenStudy (anonymous):

Nope, I do not have the answer. But can you walk me through how you did it?

OpenStudy (anonymous):

see its just the normal division

OpenStudy (anonymous):

tell me the part u don't understand

OpenStudy (anonymous):

I do not even know how to start it!

OpenStudy (anonymous):

k...lets make it simple show me some normal division..say fr example 27 divided by 2

OpenStudy (anonymous):

See if I can do it on a calculator, I can give you the answer. 13.5 I also do math weird too. Lets see if i was doing it by hand I would start off finding half of 27, which is 13.5.

OpenStudy (anonymous):

no..not that way ...i'm showing it to u

OpenStudy (anonymous):

2| 27 |13 -2 ------ 07 -06 ----- 1 this is the way i've done with ur question

OpenStudy (anonymous):

see in this type of division if u multiply 2 * 13 and add 1 u get 27

OpenStudy (anonymous):

here 1 is the remainder ..... 13 quotient ........and 27 dividend ...2 divisor

OpenStudy (anonymous):

was this much helpful?

OpenStudy (anonymous):

Yes, it was, i can divide evens good, its the odds i dont do well, lol. So ok, how would we start off this problem: y^2+8 /y^4 +2y-3

OpenStudy (anonymous):

y^2 /y^4 =y^2

OpenStudy (anonymous):

y^2/2y=?=1/2

OpenStudy (anonymous):

yes thats how i worked it out multiplying the eq by y^2 gives y^4...which help me to cancel frst term

OpenStudy (anonymous):

but in eq fr example quadratic ax^2 + bx+ c=0 a,b,c are constants now what if b =0 !!! think about it

OpenStudy (anonymous):

See if I can do it on a calculator, I can give you the answer. 13.5 I also do math weird too. Lets see if i was doing it by hand I would start off finding half of 27, which is 13.5.

OpenStudy (anonymous):

it would cancel the b out, right, and the x because it goes with the b.

OpenStudy (anonymous):

well it would be zero...........and most imprtnt thng is u keep dividing the term as long as the first time cancels out the rest simply adds up into the eq

OpenStudy (anonymous):

sometimes the remainder would be 0, while at some other time it may be some const .....like some natural number ...........and sometimes if ur unlucky there is a variable...in the remainder (like in this one ....no matter how long i would go remainder would exist in variable form)

OpenStudy (anonymous):

See if I can do it on a calculator, I can give you the answer. 13.5 I also do math weird too. Lets see if i was doing it by hand I would start off finding half of 27, which is 13.5.

OpenStudy (anonymous):

OH OK, I think I get it. I have more problems that need solving do you want to try a hand at them?

OpenStudy (anonymous):

but only a few ones( they r lengthy and boring :( )....we'l both solve it n try cmpr our answers...

OpenStudy (anonymous):

See if I can do it on a calculator, I can give you the answer. 13.5 I also do math weird too. Lets see if i was doing it by hand I would start off finding half of 27, which is 13.5.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

See if I can do it on a calculator, I can give you the answer. 13.5 I also do math weird too. Lets see if i was doing it by hand I would start off finding half of 27, which is 13.5.

OpenStudy (anonymous):

I think I might actually know how to do this one, but I didnt see in my book it worked out like this. Give me some time and I will work it out. 9a^2 -16b^2 / 3a -4b

OpenStudy (anonymous):

its a easy one

OpenStudy (anonymous):

try coverting the eq. into a^2 -b^2 then apply the identity a^2 - b^2=(a-b)(a+b)...now u'l see denominator gets canceled

OpenStudy (anonymous):

9a^2/3a =3a -16b^2 /-4b =4b 3a+4b

OpenStudy (anonymous):

yes !!! thats perfect

OpenStudy (anonymous):

See if I can do it on a calculator, I can give you the answer. 13.5 I also do math weird too. Lets see if i was doing it by hand I would start off finding half of 27, which is 13.5.

OpenStudy (anonymous):

So then do I do this: (3a+4b)(3a-4b) then do foil?

OpenStudy (anonymous):

??? i didn't get it

OpenStudy (anonymous):

ur last reply ididn't get it?

OpenStudy (anonymous):

If I do foil I get the same thing. that equals 9a^2-16b^2

OpenStudy (anonymous):

yea ..check the question guess the easiest method and get on it

OpenStudy (anonymous):

I said so then do I do this: (3a+4b)(3a-4b) Then do FOIL?

OpenStudy (anonymous):

So what is the answer?

OpenStudy (anonymous):

3a +4b

OpenStudy (anonymous):

so r u feeling okay if u get these question?

OpenStudy (anonymous):

I understand how to do this one because it has a a and a b in it. So I can divide each one then do FOIL. the others I am so lost still. So do I do the FOIL? Because that answer was 9a^2-16b^2

OpenStudy (anonymous):

no why would u do the foil if u can have answer by much better and easy mehtod

OpenStudy (anonymous):

answer is 3a+4b not 9a^2 - 16b^2.....even if its in the book ...subjected to the question u gave me